\begin{thebibliography}{99}
	\bibitem{AFM} V. Agostiniani, M. Fogagnolo, L. Mazzieri, {\it Sharp geometric inequalities for closed hypersurfaces in manifolds with nonnegative Ricci curvature}, Invent. Math. 222 (2020), no. 3, 1033--1101.
	
	\bibitem{ANT-Talenti-a}
	A.Alvino, C. Nitsch, C. Trombetti, \emph{A Talenti comparison result for solutions to elliptic problems with Robin boundary condition}. Preprint, arXiv:1909.11950.
	
	\bibitem{ALT90}A. Alvino, P.-L. Lions, G. Trombetti, {\it 
		Comparison results for elliptic and parabolic equations via Schwarz
		symmetrization}, Ann. Inst. Henri Poincar\'{e}, Analyse non
	lin\'{e}aire, 7(1990), p37-65.
	
	\bibitem{AGM} V. Amato,  A. Gentile, A.L. Masiello,	{\it Comparison results for solutions to $p$‑Laplace equations
		with Robin boundary conditions}, Ann. Mat. Pura Appl.(2021), doi.org/10.1007/s10231-021-01153-y.
	
	\bibitem{Baernstein19}A. Baernstein, Symmetrization in analysis. With David Drasin and Richard S. Laugesen. With a foreword by Walter Hayman. New Mathematical Monographs, 36. Cambridge University Press, Cambridge (2019).
	

	\bibitem{BK202012}Z. M. Balogh, A. Krist\'{a}ly, {\it Sharp geometric inequalities in spaces with nonnegative Ricci curvature and Euclidean volume growth}, arXiv:2012.11862.
	


	
	\bibitem{Bossel} M.H. Bossel, {\it  Membranes \'{l}astiquement li\'{e}s: extension du th\'{o}r\`{e}me de Rayleigh-Faber-Krahn et de
		l'in\'{g}alit\'{e} de Cheeger}, C. R. Acad. Sci. Paris S\'{e}r. I Math. 302(1), 47–50 (1986).
	
	\bibitem{Brendle} S. Brendle, {\it Sobolev inequalities in manifolds with nonnegative curvature}, 	arXiv: 2009.13717.	
	
	\bibitem{BD} D. Bucur, D. Daners, {\it An alternative approach to the Faber-Krahn inequality for Robin problems},
	Calc. Var. Partial Differential Equations 37 (2010),  75--86.
	
%	\bibitem{BG10} D. Bucur, A. Giacomini,  {\it A variational approach to the isoperimetric inequality for the Robin eigenvalue problem},
%	Arch. Rational Mech. Anal. 198 (2010), 927--961.
	
%	\bibitem{BG15} D. Bucur, A. Giacomini, {\it Faber-Krahn inequalities for the Robin-Laplacian: A free discontinuity approach},
%	Arch. Rational Mech. Anal. 218 (2015),  757--824.
	
	\bibitem{BG15} D. Bucur, A. Giacomini.
	\emph{The Saint-Venant inequality for the Laplace operator with Robin boundary conditions},
 Milan J. Math., 83 , no. 2, 327-343, 2015.
	
	
	%\bibitem{BG15AIHP} D. Bucur, A. Giacomini, {\it  Shape optimization problems with Robin conditions on the free boundary}, Ann. Inst. H. Poincaré Anal. Non Linéaire 33 (2016), no. 6, 1539–1568.
	
	
	\bibitem{Chavel} I. Chavel, Eigenvalues in Riemannian Geometry, Academic Press, Orlando, 1984.
	
\bibitem{CCL2021} D.G. Chen, Q.-M, Cheng, H. Li,	{\it Faber-Krahn inequalities for Robin Laplacian on bounded domain  in Riemannian manifolds}, Preprint. 

\bibitem{CL2021} D.G. Chen, H. Li, {\it
	Talenti's comparison theorem for Poisson equation and applications on Riemannian manifold with nonnegative Ricci curvature},	arXiv:2104.05568 

\bibitem{CGNT2021}F. Chiacchio, N. Gavitone, C. Nitsch, C.Trombetti, {\it 
Sharp estimates for the Gaussian torsional rigidity with Robin boundary conditions}, arXiv:2109.10117.




\bibitem{ColladayLM18} D. Colladay, J. Langford, and P. McDonald, {\it Comparison Results, Exit Time
	Moments, and Eigenvalues on Riemannian Manifolds with a Lower Ricci Curvature Bound}, J. Geom. Anal. 28 (2018), 3906–3927
		
		\bibitem{DaiFu2011}Q.Y. Dai, Y. X. Fu, {\it  Faber-Krahn inequality for Robin problems involving $p$-Laplacian}, Acta Math. Appl. Sin. Engl. Ser. 27 (2011), no. 1, 13-28.
	
	
	\bibitem{D1} D. Daners, {\it
		A Faber-Krahn inequality for Robin problems in any space dimension}, Math. Ann.  333 (2006), 767--785.
		
%	\bibitem{DFK17}  D. Bucur, P. Freitas, J. B. Kennedy, {\it The Robin problem},  In: A. Henrot (Ed.), Shape optimization
%	and spectral theory (De Gruyter, 2017), pp. 78–119.

\bibitem{FM202012}M. Fogagnolo and L. Mazzieri, {\it Minimising hulls, p-capacity and isoperimetric inequality on complete Riemannian manifolds}, arXiv: 2012.09490.

\bibitem{GHM2015}N. Gamara, A. Hasnaoui, and A. Makni, {\it Torsional rigidity on compact Riemannian manifolds with lower Ricci curvature bounds}, Open Mathematics 13 (2015),  557–570.
		
	\bibitem{Gromov}M. Gromov, Metric structures for Riemannian and non-Riemannian spaces. Based on the 1981 French original. With appendices by M. Katz, P. Pansu and S. Semmes. Translated from the French by Sean Michael Bates. Reprint of the 2001 English edition. Modern Birkh\"{a}user Classics. Birkhäuser Boston, Inc., Boston, MA, 2007.
	
	\bibitem{Kawohl} B. Kawohl. Rearrangements and convexity of level sets in PDE. Lecture Notes in Mathematics, 1150. Springer-Verlag, Berlin, 1985. 
		
\bibitem{Kesavan88} S. Kesavan,  {\it Some remarks on a  result of Talenti}, Annali Sc. Norm Sup. Pisa. Cl. Sci. (4) 15 (1988), 453--465.	

	\bibitem{Kesavan2006} S. Kesavan.  Symmetrization \& applications. Series in Analysis, 3, World Scientific Publishing Co. Pte. Ltd., Hackensack, NJ, 2006.
	
	
	
%	\bibitem{Kr1} E. Krahn, {\it \"Uber eine von Rayleigh formulierte Minimaleigenschaft des Kreises}, Math. Ann.
%	94 (1925), 97--100.
	
	\bibitem{Mondin0Vedovato}A. Mondino, M. Vedovato, {\it 
		A Talenti-type comparison theorem for $\operatorname{RCD(K,N)}$ spaces and applications}, arXiv:2009.03189.
	
	\bibitem{Talenti76}
	G. Talenti, \emph{Elliptic equations and rearrangements}. Ann. Scuola Norm. Sup. Pisa Cl. Sci (4) 3 (1976).
	
	\bibitem{Talenti79}
	G. Talenti, \emph{Nonlinear elliptic equations, rearrangements of functions and Orlicz spaces}. Ann. Mat. Pura Appl. 120 (1979).
	
	\bibitem{Talenti94}
	G. Talenti, \emph{Inequalities in rearrangement invariant function spaces}. Nonlinear Analysis, Function Spaces and Applications (1994).
		
\end{thebibliography}

\documentclass[12pt]{amsart}

\usepackage[english]{babel}
\usepackage[utf8]{inputenc}
\usepackage{times}
\usepackage[T1]{fontenc}
\usepackage{graphicx}
\usepackage{amscd,amsmath,amsfonts,amstext,amssymb,amsbsy,amsopn,amsthm,eucal}
\usepackage{txfonts}
\usepackage[figuresright]{rotating}
\usepackage{longtable,booktabs}
\usepackage{rotating,booktabs}
\usepackage{caption}
\usepackage{multirow}
\usepackage{tensor}
\usepackage[table]{xcolor}
\usepackage{fancyhdr}
\usepackage{color} 
\usepackage{hyperref}

\hypersetup{
	pdftitle={Talenti-Robin},
	pdfauthor={CW},
	pdfsubject={Eigenvalue},
	pdfkeywords={Bossel-Daners, Isoperimetric inequality, Robin Laplacian, First eigenvalue}
	pdfpagelayout=SinglePage,
	pdfpagemode=UseOutlines,  pdfpagemode=UseOutlines,
	colorlinks,
	linkcolor=[rgb]{0,0,0.7},
	urlcolor=[rgb]{0,0,0.4},
	citecolor=[rgb]{0.4,0.1,0}
}

\newcommand{\R}{\mathbb{R}^n}
\newcommand{\Sp}{\mathbb{S}^n}

\newcommand{\abs}[1]{\left|#1\right|}

\newcommand{\ak}{\alpha_\kappa}
\newcommand{\akz}{\alpha_0}

\newcommand{\n}{\nabla_g}

\newcommand{\ld}{\lambda_{1,\beta}(\Omega)}
\newcommand{\ldb}{\lambda_{1,\beta}(\Omega^\sharp)}
\newcommand{\snk}{\operatorname{sn_\kappa}}

\DeclareMathOperator{\Ric}{Ric}

\def\dfrac{\displaystyle\frac}

\newtheorem{prop}{Proposition}[section]
\newtheorem{thm}[prop]{Theorem}

\newtheorem{lem}[prop]{Lemma}
\newtheorem{rem}[prop]{Remark}
\newtheorem{cor}[prop]{Corollary}
\newtheorem{defn}[prop]{Definition}

\numberwithin{equation}{section}

\begin{document}
	
	
\baselineskip=17pt
	
	
\title[Talenti comparison Robin]{ Comparison results for solutions of Poisson equations with Robin boundary on complete Riemannian manifolds}

\author{Daguang Chen,Haizhong Li, Yilun Wei}

\thanks{The authors were supported by NSFC-FWO 11961131001 and  NSFC grant No. 11831005.}
	
\subjclass[2010]{{53C44}, {53C42}}
\keywords{Bossel-Daners inequality, Isoperimetric inequality, Talenti's comparison, Lorentz norm}

\maketitle

	

\begin{abstract}	
	 In this paper, by using Schwarz rearrangement and isoperimetric inequalities, we prove comparison results for the solutions of Poisson equations on complete Riemannian manifolds with  $Ric\geq (n-1)\kappa$, $\, \kappa\geq 0$, which extends the results in \cite{ANT-Talenti-a}. Furthermore, as applications of our comparison results, we obtain the Saint-Venant inequality and Bossel-Daners inequality for Robin Laplacian.  
\end{abstract}
	
\section{Introduction}
Let $(M,g)$ be an $n$-dimensional complete Riemannian manifold with $\Ric(g)\geq (n-1)\kappa$, where $\kappa\geq 0$. For $\kappa=0$, we further assume that $M$ is noncompact with positive asymptotic volume ratio, i.e.  
$$
\operatorname{AVR}(g)=\lim\limits_{r\rightarrow\infty}\frac{|B_r(x_0)|_{g}}{\omega_n r^n }>0,
$$ 
where $B_r(x_0)$ denotes the geodesic ball in $M$ centered at $x_0$ with radius $r$, while $\omega_n$ denotes the volume of $n$-dimensional Euclidean unit ball. Let $\Omega$ be a bounded domain with smooth boundary $\partial\Omega$ on $M$.  For a positive number $\beta$ and a nonnegative
function $f$ in $L^2(\Omega)$, we consider the following problem
\begin{equation}\label{eq1}
	\left\{
	\begin{array}{ll}
		-\Delta_{g} u= f,  & \text{in}\quad  \Omega,\\\\
		\dfrac{\partial u}{\partial N} +\beta u =0, & \text{on}\quad \partial\Omega,
	\end{array}
	\right.
\end{equation}
where $N$ denotes the outer unit normal to $\partial\Omega$. For $\beta=\infty$, the equation \eqref{eq1} can be seen as  the problem equipped  with  Dirichlet boundary condition.

According to isoperimetric inequalities in \cite{Brendle} for $\kappa=0$ (See also \cite{AFM,BK202012}) and in \cite{Gromov} for $\kappa=1$, we define   
\begin{equation}\label{iso-cont}
	\ak=
	\begin{cases}
		\operatorname{AVR}(g), & \kappa=0,\\
		\frac{\abs{M}_{g}}{\abs{\mathbb{S}^n}}, & \kappa=1,
	\end{cases}
\end{equation}
where $\abs{M}_g$, $\abs{\Sp}$ denote the volume of $M$ and $\Sp$, respectively. Let $(\R(\kappa),g_\kappa)$ be Euclidean space when $\kappa=0$ and be unit sphere when $\kappa=1$ with canonical metrics. In this article, we intend to establish a comparison principle with the solution of the following problem
\begin{equation}\label{eq2}
	\left\{
	\begin{array}{ll}
		-\Delta_{g_\kappa} v=f^\sharp,  & \text{in}\quad \Omega^\sharp,\\\\
		\frac{\partial v}{\partial N}+\beta v=0, & \text{on}\quad \partial\Omega^\sharp,
	\end{array}
	\right.
\end{equation}
where $\Omega^\sharp$ is a geodesic ball on $\R(\kappa)$ satisfying $|\Omega|_{g}=\ak|\Omega^\sharp|_{g_\kappa}$ and $f^\sharp$ is the Schwarz rearrangement of $f$.

 For Poisson equations with Dirichlet boundary conditions in Euclidean spaces,  Talenti \cite{Talenti76} gave pointwise comparisons of $u^\sharp$ and $v$. Talenti's comparison results were generalized to semilinear and nonlinear elliptic equations, for instance, in \cite{Talenti79, ALT90, ColladayLM18,Mondin0Vedovato}. Recently,  Talenti's comparison results were extended to  solutions of Poisson equations on complete noncompact Riemannian manifolds with nonnegative Ricci curvature by  the first two authors in \cite{CL2021}.  We also refer the reader to excellent books \cite{Baernstein19,Kawohl,Kesavan2006} for related topics.

 Recently, Alvino et al. \cite{ANT-Talenti-a} obtained the Talenti type comparison results of Poisson equations with Robin boundary conditions in Euclidean spaces.   The results in \cite{ANT-Talenti-a} were  generalized to the case of $p$-Laplace 
 operator with Robin boundary conditions in \cite{AGM} and to the torsion problem 
 for the Hermite operator with Robin boundary 
 conditions in \cite{CGNT2021}. Following the strategies in \cite{ANT-Talenti-a}, we focus on Poisson equations with Robin boundary conditions on complete Riemannian manifolds  and prove the following comparison results.
\begin{thm}[]\label{thm:main1}
	Let $(M,g)$ be an $n$-dimensional complete Riemannian manifold with $\Ric(g)\geq (n-1)\kappa$, where $\kappa=0$ or $1$. For $\kappa=0$, we further assume that $M$ is noncompact with $\operatorname{AVR}(g)>0$. Let $\Omega$ be a bounded domain with smooth boundary on $M$ and $\Omega^\sharp$ be a geodesic ball on $\R(\kappa)$ satisfying $|\Omega|_{g}=\ak|\Omega^\sharp|_{g_\kappa}$, where $\ak$ is defined in \eqref{iso-cont}. Let $u$ and $v$ be solutions to \eqref{eq1} and \eqref{eq2} respectively, then
\begin{equation}\label{thm:main1-1}
		||u||_{L^{p,1}(\Omega)}\leq \ak^{\frac{1}{p}}||v||_{L^{p,1}(\Omega^\sharp)}\quad \text{for} ~ \kappa=0,1 ~ \text{and} ~ 0< p\leq\frac{n}{2n-2},
	\end{equation}
\begin{equation}\label{thm:main1-2}
		||u||_{L^{2p,2}(\Omega)}\leq \ak^{\frac{1}{2p}}||v||_{L^{2p,2}(\Omega^\sharp)}\quad \text{for} ~ \kappa=0 ~ \text{and} ~  0< p\leq \frac{n}{3n-4},
	\end{equation}
\begin{equation}\label{thm:main1-4}
	||u||_{L^{2p,2}(\Omega)}\leq \ak^{\frac{1}{2p}}||v||_{L^{2p,2}(\Omega^\sharp)}\quad \text{for} ~ \kappa=1 ~ \text{and} ~ 0< p\leq \frac{n}{3n-3}.
\end{equation}
In particular, when $n=2$, $\kappa=1$ and $0< p\leq 1$,
\begin{equation}\label{thm:main1-5}
	||u||_{L^{2p,2}(\Omega)}\leq \ak^{\frac{1}{2p}}||v||_{L^{2p,2}(\Omega^\sharp)}.
\end{equation}
\end{thm}

\begin{thm}[]\label{thm:main2}
Under the same assumptions as Theorem \ref{thm:main1}, letting  $u$ and $v$ be solutions to \eqref{eq1} and \eqref{eq2} for $f\equiv 1$, respectively, we have
\begin{equation}\label{thm:main2-1}
	||u||_{L^{p,1}(\Omega)}\leq \ak^{\frac{1}{p}}||v||_{L^{p,1}(\Omega^\sharp)}\quad \text{for} ~ \kappa=0,1 ~ \text{and} ~ p\leq\frac{n}{n-2},
\end{equation}
and
\begin{equation}\label{thm:main2-2}
	||u||_{L^{2p,2}(\Omega)}\leq \ak^{\frac{1}{2p}}||v||_{L^{2p,2}(\Omega^\sharp)}\quad \text{for} ~ \kappa=0 ~ \text{and} ~ 0< p\leq\frac{n}{n-2}.
\end{equation}
In particular, for $n=2$ and  $\kappa=0$ , we have a pointwise comparison result
\begin{equation}\label{est-point}
	u^\sharp(x)\leq v(x) \quad \text{for all} ~ x\in \Omega^\sharp.
\end{equation}
\end{thm}

Let $(M, g)$ be a complete Riemannian manifold and let $\Omega\subset M$ be a smoothly bounded  domain in $M$.  Assume that $u$ is a positive solution to the following  torsion problem with Robin boundary condition
\begin{equation}\label{equ:tor}
	\left\{
	\begin{array}{ll}
		-\Delta_g u= 1,  & \text{in $\Omega$},\\\\
		\dfrac{\partial u}{\partial N} +\beta u =0, & \text{on $\partial\Omega$}.
	\end{array}
	\right.
\end{equation}
The  torsional rigidity $T_\beta(\Omega)$ with Robin boundary condition  of the domain $\Omega$ is defined by
\[
T_\beta(\Omega) = \int_{\Omega} u\, dV_g.
\]

In 2015, by using free discontinuity techniques, Bucur et al. in \cite{BG15} proved the Saint-Venant inequality for Robin Laplacian in Euclidean spaces. In 2019,  Alvino et al. in \cite{ANT-Talenti-a} obtained the same Saint-Venant inequality for Robin Laplacian via a Talenti type comparison result in Euclidean spaces. 

Choosing $p=1$ in  \eqref{thm:main2-1}, we  can deduce the Saint-Venant inequality for Robin Laplacian on given manifolds.
\begin{cor}[]\label{cor:tor}
Under the same assumptions as Theorem \ref{thm:main1}, for the torsion problem \eqref{equ:tor},  we obtain
	\begin{equation}\label{cor:tor-1}
		T_\beta(\Omega) \leq \ak T_\beta(\Omega^\sharp).
	\end{equation}
\end{cor}
\begin{rem}
	For the torsion problem with Dirichlet boundary condition on  Riemannian manifold $(M,g)$ satisfying $\Ric(g)\geq (n-1)$, the inequality \eqref{cor:tor-1} is due to \cite{ColladayLM18} and \cite{GHM2015}. In \cite{CL2021}, the first two  authors obtained  sharp estimates for the $L^1$-moment spectrum and  $L^\infty$-moment spectrum with Dirichlet boundary conditions on bounded domains in complete Riemannian manifolds with nonnegative Ricci curvature.
\end{rem}


 Let $\Omega$ be a bounded domain with smooth boundary on $M$ and $\Omega^\sharp$ be a geodesic ball on $\R(\kappa)$ satisfying $|\Omega|_{g}=\ak|\Omega^\sharp|_{g_\kappa}$. Let $\ld$ and $\ldb$ be the first eigenvalues to \eqref{eig-1} and \eqref{eig-2} respectively, i.e.
\begin{equation}\label{eig-1}
	\left\{
	\begin{array}{ll}
		-\Delta_{g} u= \ld u,  & \text{in $\Omega$},\\\\
		\dfrac{\partial u}{\partial N} +\beta u =0, & \text{on $\partial\Omega$},
	\end{array}
	\right.
\end{equation}
and
\begin{equation}\label{eig-2}
	\left\{
	\begin{array}{ll}
		-\Delta_{g_\kappa} v=\ldb v,  & \text{in $\Omega^\sharp$},\\\\
		\dfrac{\partial v}{\partial N}+\beta v=0, & \text{on $\partial\Omega^\sharp$}.
	\end{array}
	\right.
\end{equation}
As another application to Theorem \ref{thm:main1}, we use Talenti comparisons to deal with Bossel-Daners inequality for $n=2$. When $n>2$, however, $p=1$ no longer satisfies the conditions of \eqref{thm:main1-2} and \eqref{thm:main1-5}, hence we fail to use Theorem \ref{thm:main1} to give a direct proof. 
\begin{thm}[]\label{thm:eig}
	Let $(M,g)$ be an $n$-dimensional complete Riemannian manifold with $\Ric(g)\geq (n-1)\kappa$, where $\kappa=0$ or $1$. For $\kappa=0$, we further assume that $M$ is noncompact with $\operatorname{AVR}(g)>0$. Let $\Omega$ be a bounded domain with smooth boundary on $M$ and $\Omega^\sharp$ be geodesic ball on $\R(\kappa)$ satisfying $|\Omega|_{g}=\operatorname{\alpha_\kappa}|\Omega^\sharp|_{g_\kappa}$. Then
	\begin{equation}\label{thm:eig-1}
	\lambda_{1,\beta}(\Omega)\geq \lambda_{1,\beta}(\Omega^\sharp).
\end{equation}	
Moreover, the inequality holds in \eqref{thm:eig-1} if and only if $(M,g)$ is isometric to $(\R(\kappa),g_\kappa)$ and $\Omega$ is isometric to a geodesic ball in $(\R(\kappa),g_\kappa)$.
\end{thm}

\begin{rem}The inequality \eqref{thm:eig-1} is also called Faber-Krahn inequality  for Robin Laplacian.  When $(M,g)$ is Euclidean space, the result is due to Bossel in \cite{Bossel} for $n=2$ and Daners in \cite{D1} for all dimensions (See also \cite{DaiFu2011,BD,BG15}). For bounded domains in Riemannian manifolds with $\Ric(g)\geq n-1$, the inequality \eqref{thm:eig-1} is due to Chen-Cheng-Li in \cite{CCL2021}.
\end{rem}

The paper is organized as following. In Section \ref{Pre}, we recall the isoperimetric inequality of complete Riemannian manifolds with  $Ric\geq (n-1)\kappa$, where $\kappa\geq 0$,  Schwarz rearrangement, Lorentz space and Gronwall's inequality; 
In Section \ref{lemmas}, we establish some integral inequalities for the solutions to \eqref{eq1} and \eqref{eq2} by using the isoperimetric inequality;
In Section \ref{Proof of MT}, we give the proofs of Theorem \ref{thm:main1} and Theorem  \ref{thm:main2}; In Section \ref{pf:BD}, we give the proof of Theorem \ref{thm:eig}.

\section{Preliminaries}\label{Pre}

\subsection{Isoperimetric inequalities}

Let $(M,g)$ be an $n$-dimensional complete Riemannian manifold with $\Ric(g)\geq (n-1)\kappa$, where $\kappa=0$ or 1. For $\kappa=0$, we further assume that $M$ is non-compact with $\operatorname{AVR}(g)>0$. Let $\Omega$ be a bounded domain with smooth boundary on $M$ and $\Omega^\sharp$ be a geodesic ball on $\R(\kappa)$ satisfying $|\Omega|_{g}=\ak|\Omega^\sharp|_{g_\kappa}$. According to Brendle  \cite{Brendle} (See also \cite{AFM,BK202012}) and L\'{e}vy-Gromov \cite{Gromov}, we have the isoperimetric inequality
\begin{equation}\label{isoperi}
	\abs{\partial \Omega}_{g}\geq \ak \abs{\partial \Omega^\sharp}_{g_\kappa}.
\end{equation}
The equality  holds in \eqref{isoperi} if 
and only if $(M,g)$ is isometric to $(\R(\kappa),g_\kappa)$ and $\Omega$ is isometric to a geodesic ball in $\R(\kappa)$.  

For a complete noncompact $n$-dimensional Riemannian manifold $(M,g)$  with  nonnegative Ricci curvature and positive asymptotic volume growth, we also notice that the inequality \eqref{isoperi}  is proved by Agostiniani et al. in \cite[Theorem 1.8]{AFM} for $n=3$ and then extended  to $3\leq n\leq 7$ by Fogagnolo and Mazzieri in \cite{FM202012}. Furthermore, \eqref{isoperi} and its equality case still hold in ${\sf CD} (0,N)$  metric measure spaces based on the method of optimal mass transport by Balogh and Krist\'{a}ly in \cite{BK202012}.

\subsection{Schwarz rearrangement}
Let $u$ and $v$ be solutions to \eqref{eq1} and \eqref{eq2} respectively.  For $t\geq 0$ we denote by 
\begin{equation*}
	\begin{aligned}
	&U_t=\{x \in \Omega: u(x)>t\}, \quad \partial U_t^i = \partial U_t \cap \Omega, \quad \partial U_t^e= \partial U_t \cap \partial\Omega, \quad \mu(t) = |U_t|_{g},\\
	&V_t=\{x \in \Omega^\sharp: v(x)>t\}, \quad
	\phi(t) = |V_t|_{g_\kappa}.	
	\end{aligned}	
\end{equation*}

Denoting by $u_m$ and $v_m$ the minimum of $u$ and $v$ respectively, thanks to the positiveness of $\beta$ and Robin boundary conditions, we have $u_m \geq 0$ and $v_m \geq 0$. Since $v$ is radial, positive and decreasing along the radius, we have $V_t$ coincides with $\Omega^\sharp$ when $0\leq t<v_m$.
\begin{defn}
	Letting $h: \Omega \to \mathbb{R}$ be a measurable function, the distribution function of $h$ is the function $\mu_h : [0,+\infty)\, \to [0, +\infty)$ defined by
	$$
	\mu_h(t)= |\{x \in \Omega \, :\,  |h(x)|>t\}|_{g}.
	$$
	\end{defn}
\begin{defn}	
	The decreasing rearrangement $h^*:[0,|\Omega|_{g}]\to \mathbb{R}$ is defined based on the distribution function of $h$, that is
	\begin{equation}
		h^*(s)=
		\begin{cases}
			\underset{\Omega}{\textup{esssup}}\,h, & s=0,\\
			\inf \{t:\mu_h(t)\leq s \}, & s>0.
		\end{cases}
	\end{equation}
\end{defn}
\begin{defn}\label{SchwarzS}	
	The Schwarz rearrangement $h^\sharp:\Omega^\sharp\to \mathbb{R}$ is defined based on the decreasing rearrangement of $h$, that is
	\begin{equation}
		h^\sharp(x)= h^*(\ak|B^\kappa_R|_{g_\kappa}),
	\end{equation}
where $B^\kappa_R$ is the geodesic ball in $\R(\kappa)$ with radius $R$.
	\end{defn}
The distribution function of $h$ and $h^\sharp$ satisfies
\begin{equation}\label{Eqn:RearrMean}
	\mu_{h}(t)=\ak\mu_{h^{\sharp}}(t).
\end{equation}
By definition, $h$, $h^*$ and $h^\sharp$ are equi-distributed in the sense that
\begin{equation}\label{equi}
	\displaystyle{||h||_{L^p(\Omega)}=||h^*||_{L^p(0, |\Omega|_{g})}=\ak^{\frac{1}{p}}||h^\sharp||_{L^p(\Omega^\sharp)}}.
\end{equation}
Given measurable functions $h_1,h_2$ on $\Omega$, the Hardy-Littlewood inequality holds,
\begin{equation}\label{inq:HL}
	\int_{\Omega}{|h_1(x)h_2(x)| dV_{g}} \le \int_{0}^{|\Omega|_{g}}{h_1^*(s) h_2^*(s) ds}.
\end{equation}
Choosing $h_2=\chi_{\left\lbrace |u|>t\right\rbrace}$ in \eqref{inq:HL}, one has
\begin{equation}\label{inq:re}
	\int_{|u|>t}{|h_1(x)| dV_{g}} \le \int_{0}^{\mu(t)}{h_1^*(s) ds}.
\end{equation}

By strong maximum principle, both solutions $u$ and $v$  to \eqref{eq1} and \eqref{eq2}  achieve their minimum on boundaries. Hence $u$ and $v$ are strictly positive in the interior of domains. Moreover, by isoperimetric inequality \eqref{isoperi}, we have
\begin{equation}
	\begin{split}
		\ak v_m  |\partial\Omega^\sharp|_{g_\kappa} &= \ak\int_{\partial \Omega^\sharp} v(x) \, d\mu_{g_\kappa}= \frac{\ak}{\beta}\int_{\Omega^\sharp} f^\sharp \, dV_{g_\kappa}=\frac{1}{\beta} \int_{\Omega} f \, dV_{g} \\
		& = \int_{\partial \Omega} u(x) \, d\mu_{g} \\
		&\geq u_m  |\partial\Omega|_{g} \\
		&\geq u_m \ak |\partial\Omega^\sharp|_{g_\kappa},
	\end{split}
\end{equation}
which implies that 
\begin{equation}\label{minima_eq}
	 \min_\Omega u=u_m \leq v_m=\min_{\Omega^\sharp} v .
\end{equation}
An important consequence of \eqref{minima_eq} is that
\begin{equation} \label{mf}
	\mu (t) \leq |\Omega|_{g}=\ak \phi (t) \quad \text{for} \quad 0\leq t< v_m.
\end{equation} 

\subsection{Lorentz space}

\begin{defn}\label{defn:Lorentz}
Let $0<p<+\infty$ and $0<q\le +\infty$. The Lorentz space $L^{p,q}(\Omega)$ is the space of functions such that the quantity
	\begin{equation*}
		||h||_{L^{p,q}} =
		\begin{cases}
			\displaystyle{ p^{\frac{1}{q}} \left( \int_{0}^{\infty}  t^{q} \mu_h(t)^{\frac{q}{p}}\ \frac{dt}{t}\right)^{\frac{1}{q}}}, & \text{if}~~ 0<q<\infty,\\
			\displaystyle{\sup_{t>0} \, (t^p \mu_h(t))}, & \text{if}~~ q=\infty,
		\end{cases}
	\end{equation*}
is finite.
\end{defn}

It's well known that Lorentz space coincides with $L^p$ space when $p=q$ (See \cite{Talenti94} for more details). 

\subsection{Gronwall's inequality}
Let $\xi$ be a continuously differentiable function satisfying $$\tau\xi'(\tau)\leq \xi(\tau)+C$$ for all $\tau \geq\tau_0>0$, where $C$ is a non-negative constant. For all $\tau\geq \tau_0$, we have
\begin{equation}\label{Gronwall's inequality}
	\xi'(\tau)\leq\frac{\xi(\tau_0)+C}{\tau_0}.
\end{equation}

\section{Some lemmas}\label{lemmas}
Let $(M,g)$ be an $n$-dimensional complete Riemannian manifold with $\Ric(g)\geq (n-1)\kappa$, where $\kappa=0$ or 1. Moreover, for $\kappa=0$, assume $M$ is non-compact with $\operatorname{AVR}(g)>0$. Let $\Omega$ be a bounded domain with smooth boundary on $M$ and $\Omega^\sharp$ be a geodesic ball on $\R(\kappa)$ satisfying $|\Omega|_{g}=\ak|\Omega^\sharp|_{g_\kappa}$. Define a function $G_\kappa(l)$,
\begin{equation}\label{Gk}
	G_\kappa(l)=\frac{dI_\kappa}{dr}\circ I_\kappa^{-1}(l),
\end{equation}
where 
$$I_\kappa(r)=n\omega_n\ak \int_{0}^{r}{\snk^{n-1}(s) ds},$$
and	
\begin{equation*}
	\snk(s)=
	\begin{cases}
		s, & \kappa=0,\\
		\sin s, & \kappa=1.
	\end{cases}
\end{equation*}
We then have the following lemma:

\begin{lem}
	
Let u and v be solutions to \eqref{eq1} and \eqref{eq2} respectively. For almost every $t>0$, we have
\begin{equation}\label{lem1-1}
	G_\kappa(\mu(t))^2\leq \int_0^{\mu(t)}{f^*(s) ds}\cdot \left(-\mu'(t)+\frac{1}{\beta}\int_{\partial U_t^e}{\frac{1}{u}d\mu_{g}}\right),
\end{equation}
	
\begin{equation}\label{lem1-2}
	\tilde{G}_\kappa(\phi(t))^2= \ak^{-1}\int_0^{\ak\phi(t)}{f^*(s) ds}\cdot\left(-\phi'(t)+\frac{1}{\beta}\int_{\partial V_t\cap\partial\Omega^\sharp}{\frac{1}{v}d\mu_{g_\kappa}}\right),
\end{equation}
where $\tilde{G}_\kappa(r)=\ak^{-1}G_\kappa(\ak r).$
\end{lem}

\begin{proof}
		Multiplying \eqref{eq1} by $\varphi\in H^1(\Omega)$ and integrating by parts, we have 
\begin{equation}\label{10}
	\begin{aligned}
		\int_{\Omega}{\nabla_{g} u\cdot\nabla_{g} \varphi dV_{g}}+\beta\int_{\partial\Omega}{u\varphi d\mu_{g}}
		&=\int_{\Omega}{\nabla_{g} u\cdot\nabla_{g} \varphi dV_{g}}+\int_{\partial\Omega}{(-\varphi) \frac{\partial u}{\partial N}d\mu_{g}}\\
		&=\int_{\Omega}{\left[\nabla_{g} u\cdot\nabla_{g} \varphi-div_{g}(\varphi\nabla_{g} u) \right]dV_{g}}\\
		&=\int_{\Omega}{-\varphi \Delta_{g} u dV_{g}}\\
		&=\int_{\Omega}{f\varphi dV_{g}}.
	\end{aligned}
\end{equation}
Define a test function, for $h>0$,
\begin{equation*}
	\varphi_h=
	\begin{cases}
		0, & 0<u\leq t,\\
		h,& u>t+h,\\ 
		u-t,& t<u\leq t+h.\\
	\end{cases}
\end{equation*}
Choosing $\varphi=\varphi_h$ in \eqref{10}, one has
\begin{equation}
	\begin{aligned}
		&\int_{U_t\backslash U_{t+h}}{|\nabla_{g} u|_{g}^2 dV_{g}}+\beta\int_{\partial U_{t+h}^e}{uh d\mu_{g}}+\beta\int_{\partial U_{t}^e\backslash \partial U_{t+h}^e}{u(u-t) d\mu_{g}}\\
		&=\int_{U_t\backslash U_{t+h}}{f(u-t) dV_{g}}+\int_{U_{t+h}}{fh dV_{g}}.
	\end{aligned}
\end{equation}
Dividing by $h$ and letting $h\rightarrow 0+$, we have
\begin{equation}
	-\frac{d}{dt}\left(\int_{U_t}{|\nabla_{g} u|_{g}^2} dV_{g}\right)+\beta\int_{\partial U_t^e}{u d\mu_{g}}=\int_{U_t}{f dV_{g}}.
\end{equation} 
Applying co-area formula, we then obtain
\begin{equation}\label{13}
	\int_{\partial U_t^i}{|\nabla_{g} u|_{g} d\mu_{g}}+\int_{\partial U_t^e}{\beta u d\mu_{g}}=\int_{U_t}{f dV_{g}}.
\end{equation}
Define $$A=\left\{
	\begin{array}
		{l@{\quad\quad}l}
		|\nabla_{g} u|_{g}, & \partial U_t^i, \\
		\beta u, & \partial U_t^e. \\
	\end{array}
	\right.$$
Then \eqref{13} becomes
\begin{equation}\label{14}
	\int_{\partial U_t}{A d\mu_{g}}=\int_{U_t}{f dV_{g}}.
\end{equation} 
By \eqref{isoperi}, \eqref{inq:re}, \eqref{Gk}, \eqref{14} and co-area formula, we have
\begin{equation}\label{3.10}
	\begin{aligned}
		G_\kappa(\mu(t))^2&\leq|\partial U_t|_{g}^2\\
		&\leq\int_{\partial U_t}{A d\mu_{g}}\cdot\int_{\partial U_t}{A^{-1} d\mu_{g}}\\
		&=\int_{U_t}{f dV_{g}}\cdot\left(\int_{\partial U_t^i}{\frac{1}{|\nabla_{g} u|_{g}} d\mu_{g}}+\int_{\partial U_t^e}{\frac{1}{\beta u} d\mu_{g}}\right)\\
		&=\int_{U_t}{f dV_{g}}\cdot\left(-\mu'(t)+\frac{1}{\beta}\int_{\partial U_t^e}{\frac{1}{u} d\mu_{g}}\right)\\
		&\leq\int_0^{\mu(t)}{f^*(s) ds}\cdot\left(-\mu'(t)+\frac{1}{\beta}\int_{\partial U_t^e}{\frac{1}{u} d\mu_{g}}\right).\\		
	\end{aligned}
\end{equation}
Similarly, by $\tilde{G}_\kappa(\phi(t))=|\partial V_t|_{g_\kappa}$ and $\int_{V_t}{f^\sharp dV_{g_\kappa}}=\ak^{-1}\int_{0}^{\ak\phi(t)}{f^*(s) ds}$, we prove \eqref{lem1-2}.
	\end{proof}
 
\begin{lem}
	Let u and v be solutions to \eqref{eq1} and \eqref{eq2} respectively.  For all $t>v_m$, we have
	\begin{equation}\label{lem2-1}
		\int_{0}^{t}{\tau\left(\int_{\partial U_\tau^e}{\frac{1}{u}d\mu_{g}}\right)d\tau}\leq \frac{1}{2\beta}\int_{0}^{|\Omega|_{g}}{f^*(s) ds},
	\end{equation}
\begin{equation}\label{lem2-2}
	\int_{0}^{t}{\tau\left(\int_{\partial V_\tau\cap\partial\Omega^\sharp}{\frac{1}{v}d\mu_{g_\kappa}}\right)d\tau}= \frac{\ak^{-1}}{2\beta}\int_{0}^{|\Omega|_{g}}{f^*(s) ds}.
\end{equation}
	\end{lem} 
\begin{proof}
	By Fubini's theorem, \eqref{eq1} and \eqref{equi}, we have 
\begin{equation}
	\begin{aligned}
		\int_{0}^{+\infty}{\tau\left(\int_{\partial U_\tau^e}{\frac{1}{u}d\mu_{g}}\right)d\tau}&=\int_{0}^{+\infty}{d\tau\left(\int_{\partial \Omega\cap\{u\geq\tau\}}{\frac{\tau}{u}d\mu_{g}}\right)}\\
		&=\int_{0}^{+\infty}{d\tau\left(\int_{\partial \Omega}{\chi_{\{u\geq\tau\}}\frac{\tau}{u}d\mu_{g}}\right)}\\
		&=\int_{\partial \Omega}{\left(\int_{0}^{u}{\tau d\tau}\right)\frac{1}{u}d\mu_{g}}\\
		&=\frac{1}{2}\int_{\partial \Omega}{u d\mu_{g}}\\
		&=\frac{1}{2\beta}\int_{\Omega}{f dV_{g}}\\
		&=\frac{1}{2\beta}\int_{0}^{|\Omega|_{g}}{f^*(s) ds}.
	\end{aligned}
\end{equation}
Thus, for all $t> v_m$, we obtain
	\begin{equation}
	\begin{aligned}
		\int_{0}^{t}{\tau\left(\int_{\partial U_\tau^e}{\frac{1}{u}d\mu_{g}}\right)d\tau}&\leq \int_{0}^{+\infty}{\tau\left(\int_{\partial U_\tau^e}{\frac{1}{u}d\mu_{g}}\right)d\tau}\\
		&=\frac{1}{2\beta}\int_{0}^{|\Omega|_{g}}{f^*(s) ds}.
	\end{aligned}
	\end{equation}
With the fact that$\int_{\Omega^\sharp}{f^\sharp dV_{g_\kappa}}=\ak^{-1}\int_{0}^{|\Omega|_{g}}{f^*(s) ds}$ and $\partial V_t\cap \partial\Omega^\sharp=\emptyset$ when $t> v_m$, we deduce
\begin{equation}
	\begin{aligned}
		\int_{0}^{t}{\tau\left(\int_{\partial V_\tau\cap \partial\Omega^\sharp}{\frac{1}{v}d\mu_{g_\kappa}}\right)d\tau}&= \int_{0}^{+\infty}{\tau\left(\int_{\partial V_\tau\cap \partial\Omega^\sharp}{\frac{1}{v}d\mu_{g_\kappa}}\right)d\tau}\\
		&=\frac{\ak^{-1}}{2\beta}\int_{0}^{|\Omega|_{g}}{f^*(s) ds}.
	\end{aligned}
\end{equation}
\end{proof}

\section{Proofs of Theorem \ref{thm:main1} and Theorem  \ref{thm:main2}}\label{Proof of MT}
In this section, we give the proof of main Theorem \ref{thm:main1} and Theorem  \ref{thm:main2} by using the lemmas in previous section.
\begin{proof}[Proof of Theorem \ref{thm:main1}]
		Multiplying \eqref{lem1-1} by $t\mu(t)^{\frac{1}{p}}G_\kappa(\mu(t))^{-2}$ and integrating from $0$ to $\tau>v_m$, we have
\begin{equation}\label{52}
	\begin{aligned}
		\int_{0}^{\tau}{t\mu(t)^{\frac{1}{p}} dt}\leq& \int_{0}^{\tau}{-t\mu'(t) \mu(t)^{\frac{1}{p}} G_\kappa(\mu(t))^{-2} \int_{0}^{\mu(t)}{f^*(s) ds} dt}\\
		&+\int_{0}^{\tau}{\left[\frac{t}{\beta}\int_{\partial U_t^e}{\frac{1}{u}d\mu_{g}}\right] \mu(t)^{\frac{1}{p}} G_\kappa(\mu(t))^{-2} \int_{0}^{\mu(t)}{f^*(s) ds} dt}.
	\end{aligned}
\end{equation}
From \eqref{Gk}, it's obvious that $l^{\frac{1}{p}}G_0(l)^{-2}=n^{-2}(\omega_n\operatorname{\alpha_0})^{-\frac{2}{n}}l^{\frac{1}{p}-2+\frac{2}{n}}$ is non-decreasing when $0<p\leq \frac{n}{2n-2}$. Besides, $l^{\frac{1}{p}}G_1(l)^{-2}$ is also non-decreasing when $0<p\leq \frac{n}{2n-2}$. By definition of $G_1$ in \eqref{Gk}, we have
\begin{equation}
	\begin{aligned}
		\frac{d}{dl}\left(l^{\frac{1}{p}}G_1(l)^{-2}\right)&=\frac{1}{p}l^{\frac{1}{p}-1}\cdot \frac{G_1(l)-2plG_1'(l)}{G_1(l)^3}\\
		&=\frac{p^{-1}l^{\frac{1}{p}-1}}{(I_1'\circ I_1^{-1}(l))^4}\cdot (I_1'^2-2pI_1I_1'')\circ I_1^{-1}(l).
	\end{aligned}
\end{equation}
It is sufficient to show $k(r):=I_1'(r)^2-2pI_1(r)I_1''(r)\geq 0$ on $[0, \pi)$. Noticing that $$\cos r\cdot\int_{0}^{r}{\sin^{n-1}s\, ds}\cdot\sin^{-n}r\leq \frac{1}{n}$$ on $[0, \pi)$, one has
\begin{equation}
	\begin{aligned}
		k(r)&=\left(n\omega_n\operatorname{\alpha_1} \sin^{n-1} r\right)^2\left[1-2p(n-1)\frac{\cos r\cdot\int_{0}^{r}{\sin^{n-1}s\, ds}}{\sin^nr}\right]\geq 0		
	\end{aligned}
\end{equation}
when $0<p\leq \frac{n}{2n-2}$. By monotonicity of
$l^{\frac{1}{p}}G_\kappa(l)^{-2}$ when $0<p\leq \frac{n}{2n-2}$ and \eqref{lem2-1}, it follows
\begin{equation}\label{4.2}
	\begin{aligned}
		&\int_{0}^{\tau}{\left[\frac{t}{\beta}\int_{\partial U_t^e}{\frac{1}{u}d\mu_{g}}\right] \mu(t)^{\frac{1}{p}} G_\kappa(\mu(t))^{-2} \int_{0}^{\mu(t)}{f^*(s) ds} dt}\\
		&\leq |\Omega|_{g}^{\frac{1}{p}} G_\kappa(|\Omega|_{g})^{-2} \int_{0}^{|\Omega|_{g}}{f^*(s) ds}\cdot \int_{0}^{\tau}{\frac{t}{\beta}\int_{\partial U_t^e}{\frac{1}{u}d\mu_{g}} dt}\\
		&\leq \frac{|\Omega|_{g}^{\frac{1}{p}} G_\kappa(|\Omega|_{g})^{-2}}{2\beta^2}\left(\int_{0}^{|\Omega|_{g}}{f^*(s) ds}\right)^2.
	\end{aligned}
\end{equation}
By \eqref{52} and \eqref{4.2}, we have
\begin{equation}\label{54}
	\begin{aligned}
		\int_{0}^{\tau}{t\mu(t)^{\frac{1}{p}} dt}\leq& -\int_{0}^{\tau}{t\left(\mu'(t) \mu(t)^{\frac{1}{p}} G_\kappa(\mu(t))^{-2} \int_{0}^{\mu(t)}{f^*(s) ds} \right) dt}\\
		&+\frac{|\Omega_{g}|^{\frac{1}{p}} G_\kappa(|\Omega|_{g})^{-2}}{2\beta^2}\left(\int_{0}^{|\Omega|_{g}}{f^*(s) ds}\right)^2\\
		=&-\int_{0}^{\tau}{tdF_\kappa(\mu(t))}+\frac{|\Omega|_{g}^{\frac{1}{p}} G_\kappa(|\Omega|_{g})^{-2}}{2\beta^2}\left(\int_{0}^{|\Omega|_{g}}{f^*(s) ds}\right)^2,
	\end{aligned}
\end{equation}	
where $F_\kappa(l)=\int_{0}^{l}{w^{\frac{1}{p}}G_\kappa(w)^{-2}\int_{0}^{w}{f^*(s) ds} dw}$.
Integrating by parts and applying \eqref{Gronwall's inequality} , we obtain
\begin{equation}\label{55}
	\begin{aligned}
		\int_{0}^{\tau}{\mu(t)^{\frac{1}{p}} dt}+F_\kappa(\mu(\tau))\leq\frac{1}{v_m}&\Bigg[\int_{0}^{v_m}{dt \int_{0}^{t}{\mu(r)^{\frac{1}{p}} dr} }+\int_{0}^{v_m}{F_\kappa(\mu(t)) dt}\\&+\frac{|\Omega|_{g}^{\frac{1}{p}} G_\kappa(|\Omega|_{g})^{-2}}{2\beta^2}\left(\int_{0}^{|\Omega|_{g}}{f^*(s) ds}\right)^2\Bigg].
	\end{aligned}
\end{equation}
Making similar computations and using \eqref{lem1-2} and \eqref{lem2-2}, we have
\begin{equation}\label{56}
	\begin{aligned}
		\int_{0}^{\tau}{t\phi(t)^{\frac{1}{p}} dt}=& -\int_{0}^{\tau}{t\left(\operatorname{\alpha_\kappa}^{-1}\phi'(t) \phi(t)^{\frac{1}{p}} \tilde{G}_\kappa(\phi(t))^{-2} \int_{0}^{\operatorname{\alpha_\kappa}\phi(t)}{f^*(s) ds} \right) dt}\\
		&+\operatorname{\alpha_\kappa}^{-1}\int_{0}^{\tau}{\left[\frac{t}{\beta}\int_{\partial V_t\cap \partial \Omega^\sharp}{\frac{1}{v}d\mu_{g_\kappa}}\right] \phi(t)^{\frac{1}{p}} \tilde{G}_\kappa(\phi(t))^{-2} \int_{0}^{\operatorname{\alpha_\kappa}\phi(t)}{f^*(s) ds} dt}\\
		=&-\int_{0}^{\tau}{td\left(\operatorname{\alpha_\kappa}^{-\frac{1}{p}}F_\kappa(\operatorname{\alpha_\kappa} \phi(t))\right)}+\operatorname{\alpha_\kappa}^{-\frac{1}{p}}\frac{|\Omega|_{g}^{\frac{1}{p}} G_\kappa(|\Omega|_{g})^{-2}}{2\beta^2}\left(\int_{0}^{|\Omega|_{g}}{f^*(s) ds}\right)^2.
	\end{aligned}	
\end{equation}
Thus,
\begin{equation}\label{57}
	\begin{aligned}
		\int_{0}^{\tau}{(\operatorname{\alpha_\kappa}\phi(t))^{\frac{1}{p}} dt}+F_\kappa(\operatorname{\alpha_\kappa}\phi(\tau))=\frac{1}{v_m}&\Bigg[\int_{0}^{v_m}{dt \int_{0}^{t}{(\operatorname{\alpha_\kappa}\phi(r))^{\frac{1}{p}} dr} }+\int_{0}^{v_m}{F_\kappa(\operatorname{\alpha_\kappa}\phi(t)) dt}\\
		&+\frac{|\Omega|_{g}^{\frac{1}{p}} G_\kappa(|\Omega|_{g})^{-2}}{2\beta^2}\left(\int_{0}^{|\Omega|_{g}}{f^*(s) ds}\right)^2\Bigg].
	\end{aligned}
\end{equation}
Since \eqref{mf} holds, by \eqref{55} and \eqref{57}, we conclude
\begin{equation}
	\int_{0}^{\tau}{\mu(t)^{\frac{1}{p}} dt}+F_\kappa(\mu(\tau))\leq \int_{0}^{\tau}{(\operatorname{\alpha_\kappa}\phi(t))^{\frac{1}{p}} dt}+F_\kappa(\operatorname{\alpha_\kappa}\phi(\tau)).
\end{equation}
Letting $\tau\rightarrow +\infty$, we prove \eqref{thm:main1-1}.

Next, we come to prove \eqref{thm:main1-2}, \eqref{thm:main1-4} and \eqref{thm:main1-5}. 
Letting $\tau\rightarrow +\infty$ in \eqref{54} and \eqref{56}, we have
\begin{equation}\label{4.10}
		\int_{0}^{+\infty}{t\mu(t)^{\frac{1}{p}} dt}\leq\int_{0}^{+\infty}{F_\kappa(\mu(t)) dt}+\frac{|\Omega|_{g}^{\frac{1}{p}} G_\kappa(|\Omega|_{g})^{-2}}{2\beta^2}\left(\int_{0}^{|\Omega|_{g}}{f^*(s) ds}\right)^2,
\end{equation}
\begin{equation}\label{4.11}
	\int_{0}^{+\infty}{t(\operatorname{\alpha_\kappa}\phi(t))^{\frac{1}{p}} dt}=\int_{0}^{+\infty}{F_\kappa(\operatorname{\alpha_\kappa}\phi(t)) dt}+\frac{|\Omega|_{g}^{\frac{1}{p}} G_\kappa(|\Omega|_{g})^{-2}}{2\beta^2}\left(\int_{0}^{|\Omega|_{g}}{f^*(s) ds}\right)^2.
\end{equation}
We then need to prove 
$$\int_{0}^{+\infty}{F_\kappa(\mu(t)) dt}\leq \int_{0}^{+\infty}{F_\kappa(\operatorname{\alpha_\kappa}\phi(t)) dt}.$$ Multiplying \eqref{lem1-1} by $tF_\kappa(\mu(t))G_\kappa(\mu(t))^{-2}$ and integrating from $0$ to $\tau>v_m$, we get
\begin{equation}
	\begin{aligned}
		\int_{0}^{\tau}{tF_\kappa(\mu(t)) dt}\leq& -\int_{0}^{\tau}{t\left(\mu'(t) F_\kappa(\mu(t)) G_\kappa(\mu(t))^{-2} \int_{0}^{\mu(t)}{f^*(s) ds} \right) dt}\\
		&+\int_{0}^{\tau}{\left[\frac{t}{\beta}\int_{\partial U_t^e}{\frac{1}{u}d\mu_{g}}\right] F_\kappa(\mu(t)) G_\kappa(\mu(t))^{-2} \int_{0}^{\mu(t)}{f^*(s) ds} dt}.
	\end{aligned}
\end{equation}	
It's not hard to verify that $F_0(l)G_0(l)^{-2}$ is non-decreasing when $0<p\leq \frac{n}{3n-4}$. By definitions of $G_1$ and $F_1$, when $0< p\leq \frac{n}{3n-3}$, one has
\begin{equation*}
	G_1''(l)=\frac{d}{dl}\frac{n-1}{\tan(I_1^{-1}(l))}\leq 0, \quad
	G_1(l)-3plG_1'(l)\geq 0.
\end{equation*}
Hence we have
\begin{equation*}
	\begin{aligned}
		\frac{d}{dl}\left(F_1'(l)G_1(l)-2F_1(l)G_1'(l)\right)&=F_1''(l)G_1(l)-F_1'(l)G_1'(l)-2F_1(l)G_1''(l)\\
		&\geq F_1''(l)G_1(l)-F_1'(l)G_1'(l)\\
		&=\frac{1}{p}l^{\frac{1}{p}-1}G_1(l)^{-2}\left(G_1(l)-3plG_1'(l)\right)\int_{0}^{l}{f^*(t) dt}+l^{\frac{1}{p}}G_1(l)^{-1}f^*(l)\\
		&\geq 0.
	\end{aligned}	
\end{equation*}
Noticing that $F_1'(0)G_1(0)-2F_1(0)G_1'(0)=0$, we have
\begin{equation*}
	F_1'(l)G_1(l)-2F_1(l)G_1'(l)\geq 0.
\end{equation*}
Thus, $F_1(l)G_1(l)^{-2}$ is non-decreasing since
\begin{equation*}
	(F_1G_1^{-2})'(l)=G_1(l)^{-3}(F_1'(l)G_1(l)-2F_1(l)G_1'(l))\geq 0.
\end{equation*}
When $n=2$, $0<p\leq 1$, we have
\begin{equation*}
	(G_1G_1')'(l)=2\pi\operatorname{\alpha_1}\frac{d \cos(I_1^{-1}(l))}{dl}\leq 0,\quad
	G_1(l)-2plG_1'(l)\geq 0.
\end{equation*}
Hence we deduce
\begin{equation*}
	\begin{aligned}
		\frac{d}{dl}\left(F_1'(l)G_1(l)^2-2F_1(l)G_1(l)G_1'(l)\right)&=F_1''(l)G_1(l)^2-2F_1(l)(G_1G_1')'(l)\\
		&\geq F_1''(l)G_1(l)^2\\
		&=\frac{1}{p}l^{\frac{1}{p}-1}G_1(l)^{-1}\left(G_1(l)-2plG_1'(l)\right)\int_{0}^{l}{f^*(t) dt}+l^{\frac{1}{p}}f^*(l)\\
		&\geq 0.
	\end{aligned}	
\end{equation*}
Together with $F_1'(0)G_1(0)^2-2F_1(0)G_1(0)G_1'(0)=0$, we have
\begin{equation*}
	F_1'(l)G_1(l)^2-2F_1(l)G_1(l)G_1'(l)\geq 0.
\end{equation*}
Then $F_1(l)G_1(l)^{-2}$ is non-decreasing by
\begin{equation*}
	(F_1G_1^{-2})'(l)=G_1(l)^{-4}(F_1'(l)G_1(l)^2-2F_1(l)G_1(l)G_1'(l))\geq 0.
\end{equation*}
Since $F_\kappa(l)G_\kappa(l)^{-2}$ is non-decreasing, by \eqref{lem2-1} we have
\begin{equation}
	\int_{0}^{\tau}{tF_\kappa(\mu(t)) dt}\leq -\int_{0}^{\tau}{t dH_\kappa(\mu(t))}+\frac{F_\kappa(|\Omega|_{g}) G_\kappa(|\Omega|_{g})^{-2}}{2\beta^2}\left(\int_{0}^{|\Omega|_{g}}{f^*(s) ds}\right)^2,
\end{equation}
where $H_\kappa(l)=\int_{0}^{l}{F_\kappa(w)G_\kappa(w)^{-2}\int_{0}^{w}{f^*(s) ds} dw}$. Integrating by parts and applying \eqref{Gronwall's inequality},  one gets
\begin{equation}\label{4.15}
	\begin{aligned}
		\int_{0}^{\tau}{F_\kappa(\mu(t)) dt}+H_\kappa(\mu(\tau))\leq\frac{1}{v_m}&\Bigg[\int_{0}^{v_m}{dt \int_{0}^{t}{F_\kappa(\mu(r)) dr} }+\int_{0}^{v_m}{H_\kappa(\mu(t)) dt}\\&+\frac{F_\kappa(|\Omega|_{g}) G_\kappa(|\Omega|_{g})^{-2}}{2\beta^2}\left(\int_{0}^{|\Omega|_{g}}{f^*(s) ds}\right)^2\Bigg].
	\end{aligned}
\end{equation}
Analogously, by \eqref{lem1-2} and \eqref{lem2-2}, we obtain
\begin{equation}
	\begin{aligned}
		\int_{0}^{\tau}{tF_\kappa(\operatorname{\alpha_\kappa}\phi(t)) dt}=& -\int_{0}^{\tau}{t\left(\operatorname{\alpha_\kappa}^{-1}\phi'(t) F_\kappa(\operatorname{\alpha_\kappa}\phi(t)) \tilde{G}_\kappa(\phi(t))^{-2} \int_{0}^{\operatorname{\alpha_\kappa}\phi(t)}{f^*(s) ds} \right) dt}\\
		&+\operatorname{\alpha_\kappa}^{-1}\int_{0}^{\tau}{\left[\frac{t}{\beta}\int_{\partial V_t\cap \partial \Omega^\sharp}{\frac{1}{v}d\mu_{g_\kappa}}\right] F_\kappa(\operatorname{\alpha_\kappa}\phi(t)) \tilde{G}_\kappa(\phi(t))^{-2} \int_{0}^{\operatorname{\alpha_\kappa}\phi(t)}{f^*(s) ds} dt}\\
		=&-\int_{0}^{\tau}{tdH_\kappa(\operatorname{\alpha_\kappa} \phi(t))}+\frac{F_\kappa(|\Omega|_{g}) G_\kappa(|\Omega|_{g})^{-2}}{2\beta^2}\left(\int_{0}^{|\Omega|_{g}}{f^*(s) ds}\right)^2.
	\end{aligned}	
\end{equation}
Thus,
\begin{equation}\label{4.17}
	\begin{aligned}
		\int_{0}^{\tau}{F_\kappa(\operatorname{\alpha_\kappa}\phi(t)) dt}+H_\kappa(\operatorname{\alpha_\kappa}\phi(\tau))=\frac{1}{v_m}&\Bigg[\int_{0}^{v_m}{dt \int_{0}^{t}{F_\kappa(\operatorname{\alpha_\kappa}\phi(r)) dr} }+\int_{0}^{v_m}{H_\kappa(\operatorname{\alpha_\kappa}\phi(t)) dt}\\
		&+\frac{F_\kappa(|\Omega|_{g}) G_\kappa(|\Omega|_{g})^{-2}}{2\beta^2}\left(\int_{0}^{|\Omega|_{g}}{f^*(s) ds}\right)^2\Bigg].
	\end{aligned}
\end{equation}
By \eqref{mf}, \eqref{4.15} and \eqref{4.17}, we come up with
\begin{equation}
	\int_{0}^{\tau}{F_\kappa(\mu(t)) dt}+H_\kappa(\mu(\tau))\leq \int_{0}^{\tau}{F_\kappa(\operatorname{\alpha_\kappa}\phi(t)) dt}+H_\kappa(\operatorname{\alpha_\kappa}\phi(\tau)).
\end{equation}
Letting $\tau\rightarrow +\infty$ and using \eqref{4.10}, \eqref{4.11}, we prove \eqref{thm:main1-2}, \eqref{thm:main1-4} and \eqref{thm:main1-5}. 
\end{proof}

\begin{proof}[Proof of Theorem \ref{thm:main2}]
	If $f\equiv 1$, then \eqref{52} becomes
	\begin{equation}
		\begin{aligned}
			\int_{0}^{\tau}{t\mu(t)^{\frac{1}{p}} dt}\leq& \int_{0}^{\tau}{-t\mu'(t) \mu(t)^{\frac{1}{p}+1} G_\kappa(\mu(t))^{-2} dt}\\
			&+\int_{0}^{\tau}{\left[\frac{t}{\beta}\int_{\partial U_t^e}{\frac{1}{u}d\mu_{g}}\right] \mu(t)^{\frac{1}{p}+1} G_\kappa(\mu(t))^{-2} dt}.
		\end{aligned}
	\end{equation}
It's easy to show that $l^{\frac{1}{p}+1}G_\kappa(l)^{-2}$ is non-decreasing when $\kappa=0,1$ and $0<p\leq \frac{n}{n-2}$, hence
\begin{equation}\label{74}
	\int_{0}^{\tau}{t\mu(t)^{\frac{1}{p}} dt}\leq
	-\int_{0}^{\tau}{tdF_\kappa(\mu(t))}+\frac{|\Omega|_{g}^{\frac{1}{p}+2} G_\kappa(|\Omega|_{g})^{-2}}{2\beta^2}.
\end{equation}	
Following the similar arguments in proof of \eqref{thm:main1-1}, we are able to prove \eqref{thm:main2-1}. 

Noticing that $lF_0(l)G_0(l)^{-2}$ is still non-decreasing when $0<p
\leq \frac{n}{n-2}$, we adopt same  as the proof of \eqref{thm:main1-2}, \eqref{thm:main1-4} and \eqref{thm:main1-5} and finish the proof of \eqref{thm:main2-2}.

Finally, we give the proof of the pointwise comparison result \eqref{est-point}.
For $n=2$,  $\kappa=0$ and $f\equiv 1$, \eqref{lem1-1} becomes
\begin{equation}\label{new-1}
	4\pi\akz\leq -\mu'(t)+\frac{1}{\beta}\int_{\partial U_t^e}{\frac{1}{u}d\mu_{g}},
\end{equation}
while \eqref{lem1-2} yields
\begin{equation}\label{new-2}
	4\pi=-\phi'(t)+\frac{1}{\beta}\int_{\partial V_t\cap\partial\Omega^\sharp}{\frac{1}{v}d\mu_{g_\kappa}}.
\end{equation}
Multiplying \eqref{new-1} by $t$, integrating form 0 to $\tau>v_m$ and applying \eqref{lem2-1}, we obtain
\begin{equation}
	2\pi\akz\tau^2\leq \int_{0}^{\tau}{-t\mu' dt}+\frac{\abs{\Omega}_g}{2\beta^2}.
\end{equation}
Meanwhile, by \eqref{new-2}, one has
\begin{equation}
	2\pi\akz\tau^2=\int_{0}^{\tau}{-t(\akz\phi(t))' dt}+\frac{\abs{\Omega}_g}{2\beta^2}.
\end{equation}
Thus,
\begin{equation}
	\int_{0}^{\tau}{-t\mu' dt}\geq\int_{0}^{\tau}{-t(\akz\phi(t))' dt}.
\end{equation}
Integrating by parts and using \eqref{Gronwall's inequality}, for all $\tau>v_m$, we have
\begin{equation}
	\mu(\tau)-\akz\phi(\tau)\leq\frac{1}{v_m}\int_{0}^{v_m}{\mu(t)-\akz\phi(t) dt}.
\end{equation}
Together with \eqref{mf}, we find out that for all $\tau\geq 0$, 
\begin{equation}
	\mu(\tau)\leq\akz\phi(\tau).
\end{equation}
By definition of Schwarz rearrangement, we complete the proof of \eqref{est-point}.
	\end{proof}

  \section{Proof of Theorem \ref{thm:eig}}\label{pf:BD}
  
  For $n=2$, it is independently interesting to prove Bossel-Daners inequality by using our Theorem \ref{thm:main1}. For $n\geq 2$, the methods in \cite{Bossel} and \cite{D1}  can be adopted  to prove the Bossel-Daners inequality. For bounded domains in Riemannian manifolds with $\Ric(g)\geq (n-1)$, the inequality \eqref{thm:eig-1} has already been proven in \cite{CCL2021}. Here we only give the proof of \eqref{thm:eig-1} for complete noncompact Riemannian manifolds with nonnegative Ricci curvature and positive asymptotic volume growth.
  
  \begin{proof}[Proof of Theorem \ref{thm:eig}] Following the idea in \cite{Kesavan88}, we can prove the  Bossel-Daners inequality for the first eigenvalue of  Robin Laplacian in dimension $2$. 	Let $u$ and $z$ be solutions to \eqref{eig-1} and \eqref{eig-2} respectively.
When $n=2$, by choosing $p=1$ in \eqref{thm:main1-2} and \eqref{thm:main1-5}, we have
\begin{equation}
	\int_{\Omega^\sharp}{(u^\sharp)^2 dV_{g_\kappa}}=\ak^{-1}\int_{\Omega}{u^2 dV_{g}}\leq \int_{\Omega^\sharp}{z^2 dV_{g_\kappa}}.
\end{equation}
By Cauchy-Schwarz inequality, one has
\begin{equation}
	\int_{\Omega^\sharp}{u^\sharp z dV_{g_\kappa}}\leq \int_{\Omega^\sharp}{z^2 dV_{g_\kappa}}.
\end{equation}
By Rayleigh quotient for eigenvalues, we obtain
\begin{equation}
	\begin{aligned}
		\lambda_{1,\beta}(\Omega)&=\frac{\int_{\Omega^\sharp}{|\nabla_{g_\kappa} z|_{g_\kappa}^2 dV_{g_\kappa}}+\beta\int_{\partial \Omega^\sharp}{z^2 d\mu_{g_\kappa}}}{\int_{\Omega^\sharp}{u^\sharp z dV_{g_\kappa}}}\\
		&\geq \frac{\int_{\Omega^\sharp}{|\nabla_{g_\kappa} z|_{g_\kappa}^2 dV_{g_\kappa}}+\beta\int_{\partial \Omega^\sharp}{z^2 d\mu_{g_\kappa}}}{\int_{\Omega^\sharp}{z^2 dV_{g_\kappa}}}\geq \lambda_{1,\beta}(\Omega^\sharp).
	\end{aligned}	
\end{equation}	
Now we are in position to give  the  Bossel-Daners inequality for  $n\geq2$ and $\kappa=0$.
Let $B^0_R$ be the geodesic ball of radius $R$ in Euclidean space $(\R(0),g_0)$ such that $\abs{\Omega}_g=\operatorname{\alpha_0}\abs{B^0_R}_{g_0}$.  Let $u_0$ be the eigenfunction associated to  $\lambda_{1,\beta}(B^0_R)$  on geodesic ball $B^0_R$, i.e. 
\begin{equation}\label{robin-ball}
	\begin{cases}
		u_0^{\prime\prime}+\frac{n-1}{r}u_0^\prime+\lambda_{1,\beta}(B^0_R)u_0=0, \quad \text{on}~B^0_R,\\
		u_0^{\prime}(0)=0, \quad u_0^{\prime}(R)+\beta u_0(R)=0.
	\end{cases}	
\end{equation}
It is known that the function $u_0(r)$ is positive for $r\in [0,R]$.
Define  $v(r)=(\ln u_0)^\prime(r)$. According to Lemma 2.1 and Proposition 2.2  in \cite{CCL2021} (See also \cite{D1,BD}), the function $v(r)=\left(\ln u_0\right)^\prime(r)$ is strictly decreasing  and then $0\leq -v(r)<\beta$ for $r\in [0,R)$ and the first eigenvalue $\lambda_{1,\beta}(B^0_R)$ of Robin Laplacian  is strictly decreasing  with respect to radius $R$.
Let  $u$ be the first eigenfunction of Laplacian equation on $\Omega$. We can choose $u>0$ and normalize it such that
\begin{equation*}
	\begin{aligned}
		&\underset{x\in \overline{\Omega}}{\max} \, u(x)=1,\quad
		\underset{x\in \overline{\Omega}}{\min} \, u(x)=m.
	\end{aligned}
\end{equation*}
For $t \in (m,1) $,  set
\begin{equation}\label{set}
		U_t =\{x \in \Omega: \  u(x) >t\},\quad 
		\partial U_t^i=\{x\in\Omega: u(x)=t\},\quad
		\partial U_t^e=\partial U_t\cap \partial \Omega=\{x\in\partial\Omega: u(x)\geq t \}.
\end{equation}
Define the admissible subset $M _ { \beta } (\Omega)$ of $C(\Omega)$ by
\begin{equation}\label{Adm-M}
	M _ { \beta }(\Omega)  = \left\{ \varphi\in C ( \Omega ) : \,\varphi(x)\geq 0, \limsup _ { x \rightarrow z } \varphi ( x ) \leq \beta, \quad \text { for all } z \in \partial \Omega \right\}.
\end{equation}
For $\varphi\in M_\beta(\Omega)$, we can define the functional  $H_{\Omega}(U_t,\varphi)$ as in \cite{Bossel} and \cite{D1}
\begin{equation}\label{DB-functional}
	H_{\Omega}(U_t,\varphi)=\frac{1}{\abs{U_t}_g}\left(\beta \abs{\partial U_t^e}_g+\int_{\partial U_t^i}\varphi d\mu_g-\int_{U_t}\varphi^2dV_g\right).
\end{equation}
For the first eigenfunction $u$, one can infer that $\frac{\abs{\n u}_g}{u}\in  M _ { \beta }(\Omega)$ and 
\begin{equation}\label{DB-functional1}
	\ld=H_{\Omega}\left(U_t,\frac{\abs{\n u}_g}{u}\right),\qquad \text{for almost all}\, t\in (m,1).
\end{equation}
For $\varphi\in M_\beta(\Omega)$, set
\begin{equation*}
	w=\varphi-\frac{\abs{\n u}_g}{u},\qquad \text{and}\qquad F(t)=\int_{U_t}w\frac{\abs{\n u}_g}{u}dV_g.
\end{equation*}
From Lemma 3.1 in \cite{CCL2021} (See also \cite{BD,D1}), the functional $H_{\Omega}$ satisfies
\begin{equation}\label{equ3.0}
	H_{\Omega}(U_t,\varphi)=\ld-\frac{1}{\abs{U_t}_g}\left(\frac{1}{t}\frac{d}{dt}\left(t^2F(t)\right)+\int_{U_t}w^2dV_g\right)
\end{equation}
for almost all $t\in (m,1)$. Furthermore, suppose that $\varphi\neq\frac{\abs{\n u}_g}{u} $, then there exits  a set $S\subset (m,1)$ with $\abs{S}>0$ such that
\begin{equation}\label{DB-functional2}
	\ld>H_{\Omega}(U_t,\varphi),\qquad \text{for all}\quad t\in S.
\end{equation}
For  $t\in (m,1]$, we define the ball $B^0_{r(t)}$ with radius $r(t)$  such that $\abs{U_t}_g=\operatorname{\alpha_0}\abs{B^0_{r(t)}}_{g_0}$. For $x \in \partial U_t^i$ and $t\in (m,1]$, we define
\begin{equation}\label{funct-test}
	\varphi(x)=-v(r(t))=\frac{-u_0^\prime(r(t))}{u_0(r(t))}.
\end{equation}
It is easy to check that  $\varphi:\Omega \rightarrow (0, \infty)$   is a measurable function and  
\begin{equation}\label{}
	U_t=\{x \in \Omega: \  u(x) >t\}=\{x \in \Omega: \  \varphi(x)<-v(r(t))\}
\end{equation}
is open in $\Omega$ for $t\in (m,1)$.  From  the construction of $\varphi$ in \eqref{funct-test} and the monotonic decreasing property of $v(r)=\left(\ln u_0\right)^\prime(r)$, we notice that
\begin{equation*}
	0\leq \varphi(x)<\beta, \qquad\text{for}\quad  x\in \Omega,
\end{equation*}
which implies that $\varphi\in M_{\beta}(\Omega)$. Since $\abs{U_t}_g=\operatorname{\alpha_0}\abs{B^0_{r(t)}}_{g_0}$,
we have
\begin{equation}\label{Cavalieri}
	\int_{U_t}\varphi^2dV_g=\operatorname{\alpha_0} \int_{B^0_{r(t)}}(-v(r(t))^2dV_{g_0}, \quad 
	\quad \text{for}\, t\in (m,1].
\end{equation}
Moreover, by isoperimetric inequality \eqref{isoperi} and $\varphi=-v(r(t))\leq \beta$, we obtain that
\begin{equation}\label{iso1}
	\begin{aligned}
		\operatorname{\alpha_0}\abs{\partial B^0_{r(t)}}_{g_0}(-v(r(t)))\leq& (-v(r(t)))\abs{\partial U_t}_g\\
		\leq &\int_{\partial U_t^i}\varphi(x)d\mu_g+\beta\abs{\partial U_t^e}_g, \qquad \text{for} \quad t\in (m,1),
	\end{aligned}
\end{equation}
From \eqref{funct-test}  and \eqref{iso1}, we conclude
\begin{equation*}
	\begin{aligned}
		\ldb&=H_{B^0_R}(B^0_{r(t)},-v(r(t)))\\
		&=\frac{1}{\operatorname{\alpha_0}\abs{B^0_{r(t)}}_{g_0}}\left(\operatorname{\alpha_0}\int_{\partial B^0_{r(t)}} (-v(r(t)))d\mu_{g_0}-\operatorname{\alpha_0}\int_{B^0_{r(t)}} (-v(r(t)))^2dV_{g_0}\right)\\
		&\leq  \frac{1}{\abs{U_t}_g}\left(\beta\abs{\partial U_t^e}_g+\int_{\partial U_t^i}\varphi(x)d\mu_g-\int_{U_t}\varphi^2dV_g\right)\\
		&\leq  \ld,
	\end{aligned}
\end{equation*}
which completes the proof of the inequality \eqref{thm:eig-1}. When equality occurs in \eqref{thm:eig-1}, the above inequalities become equalities. Therefore, according to \cite{Brendle}, the equality holds in \eqref{thm:eig-1} if and only if $(M,g)$ is isometric to $(\R(0),g_0)$ and $\Omega$ is isometric to a ball $B^0_R$, where $g_0$ is the canonical metric of Euclidean space.

\end{proof}

\providecommand{\bysame}{\leavevmode\hbox
	to3em{\hrulefill}\thinspace}

\vspace{1cm}

\begin{flushleft}
	Daguang Chen,
	E-mail: dgchen@tsinghua.edu.cn\\
	Haizhong Li,
	E-mail:	lihz@tsinghua.edu.cn\\
	Yilun  Wei,
	E-mail:	weiyl19@mails.tsinghua.edu.cn\\
	Department of Mathematical Sciences, Tsinghua University, Beijing, 100084, P.R. China 	
	
\end{flushleft}

\end{document}